OpenStudy (anonymous):

If f(x)=-x^2-4x+3, find f(-2)

OpenStudy (ash2326):

We have f(x)=-x^2-4x+3 to find f(-2) , let's put x=-2 in f(x), we have f(-2)=-(-2)^2-4(-2)+3 let's simplify this f(-2)=-4+8+3 we get f(-2)=7